# On the kinks, the solitons and the shocks in discrete nonlinear transmission line

### v6

@@ Abstract
We studied discrete transmission lines constructed from ideal linear inductors and nonlinear capacitors (and possibly resistors). The localised travelling waves in the lossless transmission lines are the kinks and the solitons, which speeds and profiles were calculated. The localised travelling waves in the lossy transmission lines are the dissipative kinks and the shocks, which speeds and profiles were also calculated.
@@ II.1 The travelling waves
We take the voltages $V_{n}$ and the currents which pass though the inductors $I_{n}$ as the dynamical variables. The circuit equations (Kirchhoff laws) are
@@
Performing an integration we get
$U^{2}L\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}.$ (7)
One way to solve Eq. (7) is to perform the second integration to obtain
@@
where $F$ is the constant of integration.
However, in this Section we’ll use another method of solution of Eq. (7). After presenting $dQ/dZ$ as $(dQ/dV)(dV/dZ)$, we are motivated to introduce the new dependent variable
$E=\frac{dV}{dZ}.$ (9)
Considering $V$ as the independent variable, taking into account that
$\frac{d^{2}V}{dZ^{2}}$ $=\frac{1}{2}\frac{dE^{2}}{dV},$ (10a)
$\frac{d^{3}V}{dZ^{3}}$ $=\frac{1}{2}E\frac{d^{2}E^{2}}{dV^{2}}$ (10b)
and dividing both parts by the common multiplier $E$ in both parts, we may reduce the order of (7) and write it down as
@@
Also, from (10a) and (14) follows that in the vicinity of $V_{1,2}$
@@ II.2 The kinks
After finding $V(Z)$ we can from (1b) obtain the current
$I(Z)=-\frac{\left(V_{1}-V_{2}\right)}{2UL}\tanh\left[\frac{V_{1}-V_{2}}{2}\psi Z\right].$ (25)
@@ II.3 The solitons
Hence from (9) and (10a) follows that in the vicinity of $V_{3}$
@@
To make life easier to ourselves, let us restrict the consideration by the case $|\beta/\gamma|\gg|V_{1}|,|V_{3}|$. Physically the approximation means that $V_{1}$ and $V_{3}$ are close enough to each other, and we can ignore the cubic term in (20). In this case, excluding $V_{4}$ and introducing the amplitude $\phi$ instead of $\psi$ we obtain that (31) and (II.3) are modified to
@@
From (II.3) we see that the soliton exists both for positive and negative $\beta$. For positive $\beta$ we have dark soliton ($V_{3}<V_{1}$), and for negative $\beta$ – bright soliton ($V_{3}>V_{1}$).
@@
After finding $V(Z)$ we can from (1b) obtain the current
$I(Z)=\frac{V_{3}-V_{1}}{2UL}\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (35)
II.4 The quasi-continuum approximation
@@
Looking at the results of the previous Sections we understand that the quasi-continuum approximation can be justified if $\gamma(V_{1}-V_{2})^{2}/\alpha\ll 1$ and $\beta\left(V_{1}-V_{3}\right)/\alpha\ll 1$.
The small numerical coefficient before the sixth order derivative in the r.h.s. of (37) means that the small parameters shouldn’t be actually very small. With all that, one can expect that Eq. (37) will allow to study the cases of stronger nonlinearity than Eq. (7).
@@ III The dissipative kinks and shocks
$\tau_{R}=\frac{L}{R_{L}}+R_{C}C.$ (41)
In deriving (40) we assumed that both the influence of the higher order derivatives with respect to $Z$ and the dissipation are small, so we ignored the terms which which correspond to the products of these effects.
For the travelling wave we obtain from (40) after double integration the generalization of (8)
@@
where the ”potential energy” $\Pi(V)$ is defined by the equation (which took into account the boundary conditions)
@@
the point $V=V_{i}$ being the point of the stable/unstable equilibrium corresponds to the travelling wave being subsonic/supersonic with respect to the background. Stable equilibrium at the point $V=V_{2}$ (together with unstable equilibrium at the point $V=V_{2}$) reflect the well-known in the nonlinear waves theory fact: the shock speed is larger than the sound speed in the region before the shock but smaller than the sound speed in the region behind the shock whitham .
Let us return to (II.2) giving the sound and the kink/shock speeds for the cubic case. The kinks/shocks phase diagram in the space $V_{1},V_{2}$ can be obtained by comparing $U^{2}$ and $u^{2}(V_{2})$. Consider the case $\gamma>0,\beta\geq 0$. If $V_{1}$ and $V_{2}$ have the same sign, the travelling wave is always the shock. In the opposite case, the wave can be either the kink or the shock, depending upon $V_{1}$ and $V_{2}$. The boundary in the phase space between the kinks and the shocks is especially simple for $\beta=0$
@@
To study quantitatively dissipative kinks and shocks we will use the approach of the previous Section. After repeating exactly what was done there we obtain modified Eq. (11)
@@
At this stage we can repeat verbatim the part of Section II.2 leading to (22), hence the dissipative kinks and the shocks are of the same shape (given by (24)) as the kinks in the lossless line.
However, Eq. (II.2) is generalized to
$\psi^{2}$ $=6U^{2}LC_{0}\gamma,$ (50a)
$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}+\frac{4U\tau_{R}}{\psi}.$ (50b)
hence the value of $V_{2}$ (for a given $V_{1}$) depends upon $R$ (Eq. (50b) instead of (23b)) and so does the kink (the shock) speed and amplitude, in spite of the validity in all the cases of Eq. (21b).
We need to emphasise the different meaning of the condition (50b) for the kinks and the shocks. For the kinks, (50b) is the constraint on $V_{1},V_{2},R$. And it happens that the kink satisfying such constraint is described by the elementary function. For the shock, (50b) is just the condition for it being described by an elementary function. But, of course, the shock with any other values of $V_{1},V_{2},R$ also exists (provided that $V_{1},V_{2}$ satisfy the constraint (45)). Note also that by integrating (49) we can obtain Abel equation of the second kind, and (50b) is the condition of the solvability of the equation polyanin .
@@
Similar to Eq. (50b) in the case of shocks, Eq. (53b) is just the condition for the shock to be expressed by an elementary function. The shock exists for any other value of $\Delta V$.
@@
$V_{2}(Z)=V_{2}+\frac{V_{1}-V_{2}}{\left[\exp\left(\chi Z/2\right)+1\right]^{2}}.$ (54)
@@ IV Conclusions
We considered localised travelling waves of the voltage and the current in the discrete lossless transmission line with nonlinear capacitors. We have shown that these waves are of two kinds: the kinks, when the voltage has different asymptotic values at both sides of the transmission line, and the solitons (solitary waves), when the voltage has the same asymptotic value at both sides of the transmission line. We also studied lossy transmission lines. The localised travelling waves in such systems are the dissipative kinks and the shocks. We calculated the speeds and the profiles both of the dissipative kinks and of the shocks.

### v7

# Exact analytical solutions for the kinks, the solitons and the shocks in discrete nonlinear transmission line with nonlinear capacitance

@@ Abstract
We studied discrete transmission lines constructed from ideal linear inductors and nonlinear capacitors (and possibly resistors). The localised travelling waves in the lossless transmission lines are the kinks and the solitons, the localised travelling waves in the lossy transmission lines are the dissipative kinks and the shocks. The speeds and profiles of all these waves were calculated.
@@ II.1 The travelling waves
We take the voltages $V_{n}$ and the currents which pass though the inductors $I_{n-1,n}$ as the dynamical variables. The circuit equations (Kirchhoff laws) are
@@
After finding $V(Z)$ we can from (1b) obtain the current
$I(Z)=-\frac{V(Z)}{UL}+\text{const}.$ (7)
Performing an integration in (6) we can get
$\frac{1}{12}\frac{d^{3}V}{dZ^{3}}=U^{2}L\frac{dQ}{dZ}-\frac{dV}{dZ}.$ (8)
Notice that equation with the same form as (8) one would obtain after making the travelling wave ansatz in the generalised Korteweg-de Vries-modified Korteweg-de Vries (KDV-mKDV) equation.
We may perform the second integration to obtain
@@
where $F$ is the constant of integration. Introducing the new dependent variable
$E=\frac{dV}{dZ},$ (10)
considering $V$ as the independent variable and taking into account that
$\frac{d^{2}V}{dZ^{2}}=\frac{1}{2}\frac{dE^{2}}{dV},$ (11)
we may present (9) as
$\frac{1}{24}\frac{dE^{2}}{dV}=U^{2}LQ(V)-V+F.$ (12)
However, for technical reasons, it will be more convenient to return to (8) and avoid the integration of the equation. After presenting $dQ/dZ$ as $(dQ/dV)(dV/dZ)$, taking into account that
$\frac{d^{3}V}{dZ^{3}}=\frac{1}{2}E\frac{d^{2}E^{2}}{dV^{2}},$ (13)
and dividing both parts of (8) by the common multiplier $E$ in both parts, we may reduce the order of the equation and write it down as
@@
Also, from (11) and (17) follows that in the vicinity of $V_{1,2}$
@@
The main result of the Section II.1 is Eq. (14). The natural question arises: why we were not happy with Eq. (9) (which can be easily integrated in quadratures) and introduced instead of it the former equation. The detailed answer to this question will be given in Section III, but we can formulate it announce it right now. The main aim of the present paper is obtaining exact analytical solutions for the kinks and the shocks in the lossy transmission line. When the loss terms are included into the equations, (14) can be integrated analytically much easier than (9).
@@ II.3 The solitons
Hence from (10) and (11) follows that in the vicinity of $V_{3}$
@@
To make life easier to ourselves, let us restrict the consideration by the case $|\beta/\gamma|\gg|V_{1}|,|V_{3}|$. Physically the approximation means that $V_{1}$ and $V_{3}$ are close enough to each other, and we can ignore the cubic term in (23). In this case, excluding $V_{4}$ and introducing the amplitude $\phi$ instead of $\psi$ we obtain that (33) and (34) are modified to
@@
From (35) we see that the soliton exists both for positive and negative $\beta$. For positive $\beta$ we have dark soliton ($V_{3}<V_{1}$), and for negative $\beta$ – bright soliton ($V_{3}>V_{1}$).
@@
II.4 How good is the quasi-continuum approximation?
@@
Looking at the results of the previous Sections we understand that the fifth order derivative in Eq. (38) can be discarded if $\gamma(V_{1}-V_{2})^{2}/\alpha\ll 1$ and $\beta\left(V_{1}-V_{3}\right)/\alpha\ll 1$. The small numerical coefficient before the sixth order derivative in the r.h.s. of (38) means that the small parameters shouldn’t be actually very small.
With all that we have to admit that the transition from the discrete to the continuous system is an uncontrollable approximation. More specifically, the mere fact of existence of the travelling wave of the form (5) in the discrete system should be proved analytically or, at least, numerically. We hope to return to this issue in the future.
@@ III The dissipative kinks and shocks
$\tau_{R}\equiv\frac{L}{R_{L}}+R_{C}C.$ (42)
In deriving (41) we assumed that the dissipation is small, so we ignored the terms which correspond to the product of the dissipation coefficients and the higher order derivatives with respect to $Z$. For the travelling wave we obtain from (41) after integration the generalization of (8)
$\frac{1}{12}\frac{d^{3}V}{dZ^{3}}+U\tau_{R}\frac{d^{2}V}{dZ^{2}}=U^{2}L\frac{dQ}{dZ}-\frac{dV}{dZ}.$ (43)
Similar to what we said in Section II.1, notice that equation with the same form as (43) one would obtain after making the travelling wave ansatz in the generalised Korteweg-de Vries-modified Korteweg-de Vries (KDV-mKDV) equation with dissipation.
Performing the second integration, we obtain the generalization of (9)
@@
From (39) we obtain that in the present case the sum of the currents flowing through the nonlinear inductance and the ohmic resistance $R_{L}$ is given by the equation
$I(Z)=-\frac{V(Z)}{UL}-\frac{\tau_{R}}{L}\frac{dV(Z)}{dZ}+\text{const}.$ (45)
@@
where
@@
the point $V=V_{i}$ being the point of the stable/unstable equilibrium corresponds to the travelling wave being subsonic/supersonic with respect to the background. Stable equilibrium at the point $V=V_{2}$ (together with unstable equilibrium at the point $V=V_{1}$) reflect the well-known in the nonlinear waves theory fact: the shock speed is larger than the sound speed in the region before the shock but smaller than the sound speed in the region behind the shock whitham .
Let us return to (24) giving the sound and the kink/shock speeds for the cubic case. The kinks/shocks phase diagram in the space $V_{1},V_{2}$ can be obtained by comparing $U^{2}$ and $u^{2}(V_{2})$. Consider the case $\gamma>0,\beta\geq 0$. If $V_{1}$ and $V_{2}$ have the same sign, the travelling wave is always the shock. In the opposite case, the wave can be either the kink or the shock, depending upon $V_{1}$ and $V_{2}$. The boundary in the phase space between the kinks and the shocks is especially simple for $\beta=0$
@@
To study quantitatively dissipative kinks and shocks we will use the approach of the previous Section. After repeating exactly what was done there starting from Eq. (43), we obtain modified Eq. (14)
@@
Considering again the cubic case we can easily check up that (25) is still the solution of Eq. (52), with $\psi$ being still given by Eq. (26a), but Eq. (26b) is generalized to
$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}+\frac{4U\tau_{R}}{\psi}.$ (53)
Hence the dissipative kinks and the shocks are of the same shape (given by (27)) as the kinks in the lossless line.
Now we understand what all the trouble of introducing the new dependent and independent variables was for. We were able to integrate analytically Eq. (52) (while we didn’t see any way to integrate analytically Eq. (44)).
We need to emphasise the different meaning of the condition (53) for the kinks and the shocks. For the kinks, (53) is the condition of the solution existence (and the solution happens to be described by the elementary function). For the shock, (53) is just the condition for the solution to be described by an elementary function. But, of course, the shock with any other values of $V_{1},V_{2},\tau_{R}$ also exists (provided that $V_{1},V_{2}$ satisfy the constraint (48)). So the obtained shock can be called the particular shock of the first kind. Note also that by integrating (52) we can obtain Abel equation of the second kind, and (53) is the condition of the solvability of the equation polyanin .
@@
Similar to Eq. (53) in the case of shocks, Eq. (56b) is just the condition for the shock to be expressed by an elementary function. The shock exists for any other value of $V_{1}-V_{2}$. So the obtained shock can be called the particular shock of the second kind.
@@
$V(Z)=V_{2}+\frac{V_{1}-V_{2}}{\left[\exp\left(\chi Z/2\right)+1\right]^{2}}.$ (57)
@@ IV Conclusions
We considered localised travelling waves of the voltage and the current in the discrete lossless transmission line with nonlinear capacitors. We have shown that these waves are of two kinds: the kinks, when the voltage has different asymptotic values at both sides of the transmission line, and the solitons (solitary waves), when the voltage has the same asymptotic value at both sides of the transmission line. We also studied lossy transmission lines. The localised travelling waves in such systems are the dissipative kinks and the shocks.
The velocity of the kink, both in the lossless and in the lossy line, and of the shock in the lossy line is given by the same equations
$\frac{1}{U^{2}}=L\frac{Q_{2}-Q_{1}}{V_{2}-V_{1}},$ (58)
where $Q_{1},V_{1}$ and $Q_{2},V_{2}$ are the asymptotic values of the capacitor charge and the voltage on both ends of the transmission line.
To our opinion the most important equation of the paper is Eq. (52) describing the lossy transmission line
When the charge of the non-linear capacitor and the voltage on it are connected by the equation
$Q(V)=\text{const}+C_{0}\left(\alpha V+\beta V^{2}+\gamma V^{3}\right),$ (60)
the profiles of the kink, both in the lossless and in the lossy line, and of the particular shock of the first kind in the lossy line, can be easily calculated on the basis of the solution of Eq. (59)
$E(V)=\psi(V-V_{1})(V-V_{2}),$ (61)
and are given by the same equation
$V(Z)=\frac{V_{1}+V_{2}}{2}-\frac{V_{1}-V_{2}}{2}\tanh\left[\frac{V_{1}-V_{2}}{2}\psi Z\right],$ (62)
where $\psi^{2}=6U^{2}LC_{0}\gamma$, and the condition on the asymptotic values of the voltage on both ends of the transmission line $V_{1}$ and $V_{2}$ are given by Eq. (53). The profile of the particular shock of the second kind is given by Eq. (57.